How do I solve this quadratic equation by factoring? 2x^2+11x=6
Answer by jim_thompson5910(35256)

2x^2+11x=6

2x^2+11x-6=0

(x+6)(2x-1)=0

x+6=0 or 2x-1=0

x=-6 or 2x=1

x=-6 or x=1/2

So the two solutions are x=-6 or x=1/2